1. The radius of the planet mars is approximately 3,396 km and has a gravitational acceleration equal to three eighths times than that of the earth. What is the velocity of escape? Ans. 5 km/s
  2. Initially a tank holds 100 gal of a brine solution containing 20lb of salt. At t=0, fresh water is poured into the tank at the rate of 5 gal/min, while stirred mixture leaves the tank at the rate 3 gal/min. what is the amount of salt in the tank after 1 hours? How long will it take for the take to contain 10 lb of salt? Ans. 6.12 lb; 29.37 min.
1
Expert's answer
2021-11-17T15:41:21-0500

1.


"g_E=\dfrac{GM_E}{R_E^2}"

"g_{pl}=\dfrac{GM_{pl}}{R_{pl}^2}=\dfrac{3}{8}g_E"

The velocity of escape is


"v_{escape}=\sqrt{\dfrac{2GM_{pl}}{R_{pl}}}=\sqrt{\dfrac{2(\dfrac{3}{8}g_E)(R_{pl}^2)}{R_{pl}}}"

"v_{escape}=\sqrt{\dfrac{3g_ER_{pl}}{4}}"

"v_{escape}=\sqrt{\dfrac{3(9.81)(3396000)}{4}}=4998.6(m/s)"

"5\ km/s"


2.

Let "s(t) =" amount, in lb of salt at time "t." Then we have

"\dfrac{ds}{dt}="(rate of salt into tank) − (rate of salt out of tank)


"\dfrac{ds}{dt}=0-\dfrac{3s}{100+(5-3)t}"

So we get the differential equation


"\dfrac{ds}{dt}=-\dfrac{3s}{100+2t}"


"\int \dfrac{ds}{s}=-\int\dfrac{3dt}{100+2t}"

"s(t)=c_1(100+2t)^{-3\/2}"

"s(0)=c_1(100)^{-3\/2}=20"

"c_1=20000"

"s(t)=20000(100+2t)^{-3\/2}"

"s(60)=20000(100+2(60))^{-3\/2}"

"s(60)=6.13\ lb"


"s(t)=20000(100+2t)^{-3\/2}=10"

"(100+2t)^{3\/2}=2000"

"100+2t=100\sqrt[3]{4}"

"t=50(\sqrt[3]{4}-1)"

"t=29.37\ min"